12 March 2010

Almost done...

As we near the end of the campaign, the weather is starting to turn less cooperative. The day started off warm (-36) and clear, but clouds rolled in mid-day and lingered until sunset.

At PEARL, Cristen finished aligning the optics of the PEARL-GBS and then performed resolution and new FOV tests on in the instrument. Felicia took 38 MIR measurements with PARIS while Rodica took 56 MIR and 9 NIR measurements with the Bruker. She also performed cell tests on the instrument.

At 0PAL, Emily tested the temperature sensor configuration and worked with Graeme (off-site) to redesign the configuration of the polarotor channel and perform realignment. Zen continued work on the E-AERI's retrieval algorithm and completed checks to the Hot BB cables. He also made fixes to the instrument's config file writer.

After dinner, Emily, Zen, Keith and Cristen played another match of Frisbee golf. Paul and Rodica joined in for a great photo-op after working late at PEARL, just as the sun was setting. Hopefully the weather will hold out for our last day tomorrow!

Zen

Back to Daily Reports
Back to Home